OpenStudy (anonymous):

Find the equation of the tangent line to the curve y^3 + yx^2 + x^2 - 3y^2 = 0 at the point (0,3) @Mathematics

OpenStudy (amistre64):

seeing how the derivative of the equation defines the slope at any given point; all you have to recall is how to write the equation of a line

OpenStudy (amistre64):

y-Py=m(x-Px) ; where they give you P(0,3) y-3=m(x-0) ; and m=slope, or f' y-3=f'(0,3)(x-0) ; and m=slope, or f'

OpenStudy (anonymous):

did you get a slope of -8/9 for this one?

OpenStudy (anonymous):

just about to start ha

OpenStudy (anonymous):

alright lol let me know what you end up getting

OpenStudy (anonymous):

k man

OpenStudy (anonymous):

i dont get how the hint helps us for#1

OpenStudy (anonymous):

all I did was make it look like the definition of the derivative, then said it was equal to whatever the derivative was (which I got using the chain rule) and just subbed in whatever it was telling me to

OpenStudy (anonymous):

f(x)-f(a) /x-a?

OpenStudy (anonymous):

did u get 1

OpenStudy (anonymous):

-1 for the first one

OpenStudy (anonymous):

is what i put correct?

OpenStudy (anonymous):

then say that is equal to f' of e^sinx-1

OpenStudy (anonymous):

that's what I did. I used the definition you said there

OpenStudy (anonymous):

which is e^0*-1?

OpenStudy (anonymous):

yeah, when you evaluate the derivative at pi and that gives you -1 that's how I did it anyways

OpenStudy (anonymous):

for b i hope u got 1/80

OpenStudy (anonymous):

yeah that's what I got

OpenStudy (anonymous):

k i feel better

OpenStudy (anonymous):

hope I did it right lol

OpenStudy (anonymous):

k so for 3: derivative is: 3y^2*y'+y'x^2+y2x+2x-6y*y'=0

OpenStudy (anonymous):

then i got : y'=-2x-2xy /3y^2+x^2-6y

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

really? wasnt too confident now do i just sub in 0 and 3 for x and y

OpenStudy (anonymous):

put it all into the equation for the tangent line and sub in your x and y

OpenStudy (anonymous):

f(a) +f'(a)(x-a)?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

to get f(a) do i have to isolate y?

OpenStudy (anonymous):

f(a) is 3 assuming the point is on the function

OpenStudy (anonymous):

i dont see how u could get anything other than 0 for slope

OpenStudy (anonymous):

x is 0 and we have -2xy -2x in the numerator

OpenStudy (anonymous):

I missed an x in the numerator of the derivative ( I factored out the 2x) I think it's 0 then

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if we did it right then L(x) = 3 then right?

OpenStudy (anonymous):

thats what i got

OpenStudy (anonymous):

for 4 did you get the derivative is 2x

OpenStudy (anonymous):

you need to use logarithmic differentiation for that one

OpenStudy (anonymous):

k

OpenStudy (anonymous):

y(2xlnx+x)

OpenStudy (anonymous):

yeah that's it, then I put the x^(x^2) in for the y

OpenStudy (anonymous):

so its changing at e^-1/2 but how do i prove when <>0

OpenStudy (anonymous):

say its increasing

OpenStudy (anonymous):

you can either make a sketch or sub in values greater than e^-1/2 and less than e^-1/2 to check the behaviour

OpenStudy (anonymous):

for 5 i differentiated wrt to x then wrt y now what

OpenStudy (anonymous):

make the dx/dt = dy/dt

OpenStudy (anonymous):

ya so 16x(2-x)=9y(y-2)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

its not right but i went: differ. wrt to x 32x*x'+18y=0

OpenStudy (anonymous):

you need to diff wrt time

OpenStudy (anonymous):

so 32x*dx/dt+18y*dy/dt=144

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i dont know how to make the, equal

OpenStudy (anonymous):

them

OpenStudy (anonymous):

or does 144 become 0

OpenStudy (anonymous):

eh